# hyperrefSuppressing link with empty target

### v1

@@ Abstract
Methods. We use DYNAMICO to run a series of long-timescale, HD209458b-like, atmospheric models at various surface rotation rates. For models that are tidally-locked, we consider rotation rates between $1/16^{\mathrm{th}}$ and $40$ times the rotation rate of HD209458b, whilst for non-synchronous models we consider the range $1/8^{\mathrm{th}}$ to $4$ times HD209458b.
Results. We find that our synchronous models fall into one of three $\Omega$-dependent regimes: at low $\Omega$, we find that the outer atmosphere dynamics are driven by a divergent day-night wind, which drives weak vertical transport and can lead to the formation of a night-side hot-spot. At intermediate $\Omega$, we find classical hot Jupiter dynamics, whilst at high $\Omega$ we find a strong Coriolis effect that suppresses off-equator dynamics, including the jet-driving standing waves, thus also reducing vertical transport. As for non-synchronicity, when small, we find that it has little effect on the dynamics. However as it grows, we find that temporal variations prevents the formation of the persistent structures that drive large scale dynamics and transport.
@@ 1 Introduction
Thanks to the continuing efforts of the observational and instrumental astrophysics community, the parameter regime in which hot Jupiters can be found has expanded greatly (see, for example, Shporer et al. 2017; Hjorth et al. 2019; Huang et al. 2020; Schlecker et al. 2020; Dong et al. 2021; Stangret et al. 2022; Ulmer-Moll et al. 2022). In particular, there has been a significant increase in the range of orbital radii in which hot Jupiters can be observed. Simultaneously, numerical studies of hot Jupiter, and hot Brown Dwarf, atmospheres has revealed that rotation can play a significant role on both the horizontal and vertical dynamics (Showman et al. 2014; Sainsbury-Martinez et al. 2021). For example, Sainsbury-Martinez et al. (2021), found that, in addition to the link between radius inflation and surface irradiation, a link might also exist between the level of radius inflation observed and the surface rotation rate of the exoplanet in question. Specifically, they suggested that the lack of observed radius inflation for the ultra-short-orbit hot Brown Dwarf SDSS1411b could be explained by the influence that rotation has on the atmospheric dynamics.
Examples of how the Coriolis effect can influence atmospheric dynamics can be found in our own solar system. Generally, planetary atmospheric dynamics depend upon the influence of both surface irradiation and rotation, with the later effect leading to the formation of three distinct atmospheric regimes.
At slow rotation rates, e.g. Venus and Titan, atmospheric dynamics are dominated by planetary scale Hadley cells (Hadley 1735; Schneider 2006; Showman et al. 2013) which mix temperature horizontally, resulting in weak latitudinal temperature gradients. This is known as Cyclostrophic circulation, and leads to significant super-rotation of the winds, as well as significant, circumpolar, vortexes, all of which is described well by Thermal Wind Balance (Flasar et al. 2010).
@@
The strength of the meridional heat transport continues to decrease as we move into the rapid rotation regime, as seen on Jupiter, Saturn, and the other gaseous solar-system giants. Here the Coriolis effect plays a significant role at all latitudes, leading to a Geostrophic atmosphere dominated by the infamous banded jet structure (Ingersoll & Vasavada 1998) (with baroclinic instabilities driving mixing/angular-momentum-transport between bands).
@@
A similar change in dynamics is expected to occur in exoplanetary atmospheres, albeit over a significantly wider parameter space thanks to the much wider variety of planetary type/orbital radii combinations found outside of our solar system. One such combination found only outside of our solar-system is the hot Jupiter, a Jupiter-like, gaseous planet which is in a very short orbit and hence is very highly irradiated. Whilst a number of recent studies have explore how atmospheric dynamics changes with orbital radii (Showman et al. 2014; Sainsbury-Martinez et al. 2023a), these come with the proviso that the models link effects caused by changes in both surface irradiation and planetary rotation, making it difficult to explore the effects that rotation alone has on the dynamics.
@@
As such, and in order to isolate how rotation affects the atmospheric dynamics (with a particular focus on vertical heat transport, which is believed to be the primary driver of radius inflation - see Sainsbury-Martinez et al. 2019, 2021; Sainsbury-Martinez et al. 2023b), here we present a series of deep, HD209458b-like atmospheric models in which everything but the rotation rate is held constant. Specifically the rotation rate will be varied between $\Omega_{\mathrm{rot}}=0.0625\Omega_{0}$ and $\Omega_{\mathrm{rot}}=40\Omega_{0}$, where $\Omega_{0}=2.1\times 10^{-5}\mathrm{s^{-1}}$ is the canonical angular rotation rate of HD209458b, and the atmospheric models will be run until the deep atmosphere (with $P_{max}=200$\mathrm{bar}$$) has equilibrated. These values where chosen such that our models would span the rotation regimes found in our own solar system, and we will investigate cases in which the rotation is both synchronous and non-synchronous.
@@
The structure of this work is as follows: In section 2 we introduce the model we use to explore the effects of rotation on atmospheric dynamics and observables: This includes both a brief description of our 3D global circulation model (GCM) DYNAMICO, as well as a more indepth discussion of the parametrisations we use. Then, in section 3, we discuss how rotation influences the atmospheric dynamics of synchronously rotating hot Jupiters, starting with zonal-mean dynamics before moving onto the horizontal and vertical winds and heat transport, which we explore via both the helmholtz wind decomposition as well as the enthalpy flux. Together these allow us to investigate how rotation influences deep atmospheric heating, and hence observed radius inflation. We then, in section 4, perform a similar analysis for a set of hot Jupiter atmospheric models in which the orbital period and planetary rotation period are allowed to diverge (i.e. non-synchronous rotation), investigating how sensitive the standing-wave driven dynamics, and hence the vertical heat transport, are to longitudinal shifts in the surface irradiation profile. We finish, in section 5, with concluding remarks, discussing the implications of our results and suggestions for future studies, with next-generation GCMs, which will investigate the robustness of our results. In particular, we are very interested in confirming if the formation of a night-side hot-spot at slow rotation rates is just a feature of our atmospheric models, or if it is a robust feature that may have important implications for both the chemistry and observations of more slowly rotating hot Jupiters (i.e. hot Jupiters in longer orbits around massive/bright stars).
@@ 2.1 DYNAMICO
DYNAMICO is one of a number of highly computationally efficient global circulation models which has been developed to run on the next-generation of Exascale supercomputers, enabling atmospheric models of resolutions and regimes that where previously out of reach. To achieve this, DYNAMICO solves the primitive equations of meteorology (Dubos & Voitus 2014; Vallis 2006) on a spherical icosahedral grid (Dubos et al. 2015). Here we configure this horizontal-grid such that the angular resolution of each grid-cell is approximately $2.5^{\circ}$. Vertically, the levels of the pressure-grid are linearly spaced in $\log\left(P\right)$ space with a maximum pressure of $200$\mathrm{bar}$$ and a minimum pressure of $7\times 10^{-3}$\mathrm{bar}$$. This pressure range was selected such that we model both the highly irradiated outer atmosphere where irradiation always dominates over transport as well as a sufficiently large region of the deep atmosphere such that the effects of differing vertical heat transport rates are clearly visible. Note that the vertical boundaries of our model are closed and stress-free such that the only means of energy transport into and out-of the system are the numerical-dissipation (implemented as a hyperdiffusion filter with a fixed dissipation timescale) required to stabilise the system against grid-scale noise, and the Newtonian cooling relaxation scheme which we use to model the stellar irradiation (subsection 2.2).
Following the calibration procedures of Sainsbury-Martinez et al. (2019) and Sainsbury-Martinez et al. (2021), we select a numerical dissipation timescale of $\tau_{dissip}=2500\,\mathrm{s}$ in order to ensure the stability of the equatorially confined jets that develop in our most rapidly rotating models (e.g. 1(c)). These narrow jets with their high latitudinal velocity gradients also result in us setting a 90 second model time-step, which balances the need to correctly reproduce model physics, whilst also minimising the computational cost modelling equilibrium deep atmospheres.
@@ 2.2 Radiative forcing via Newtonian cooling
$\frac{\partial T\left(P,\theta,\phi\right)}{\partial t}=-\frac{T\left(P,\theta,\phi\right)-T_{eq}\left(P,\theta,\phi\right)}{\tau_{rad}\left(P\right)}\,.$ (1)
For numerical reasons, we express the equilibrium temperature profile ($T_{eq}\left(P,\theta,\phi\right)$) onto which the atmosphere is relaxed, in terms of the night-side temperature profile $T_{n}\left(P\right)$ (Iro et al. 2005) and a pressure dependent day-night temperature difference ($\Delta{T}(P)$) which has been calibrated such as to reproduce observations of the hot Jupiter HD209458b:
@@
Note that, for the majority of models discussed here, we set the deep radiative time-scale to infinity $\tau_{rad}^{d}=\infty$, thus disabling radiative forcing in the deep atmosphere where it is expected to be significantly slower than advection/transport (Sainsbury-Martinez et al. 2019).
@@ 2.3 HD209458b-like Atmospheric Models
In order to explore how rotation impacts the outer and deep atmospheric dynamics and observables of highly irradiated, gaseous, exoplanets, we have explored synchronously rotating atmospheres at 28 different rotation rates (between $\Omega_{\mathrm{rot}}=0.0625\Omega_{0}$ and $\Omega_{\mathrm{rot}}=40\Omega_{0}$ discussed in section 3) and non-synchronously rotating atmospheres at 13 different planetary rotation rates (between $\Omega_{\mathrm{rot}}=0.25\Omega_{0}$ and $\Omega_{\mathrm{rot}}=4\Omega_{0}$, with the orbital period held constant at that of HD209458b - discussed in section 4).
@@
Figure 1: Zonally and temporally averaged zonal wind profiles for three HD209458b-like models with rotation rates of; $0.125\Omega_{0}$ top, $\Omega_{0}$ middle, and $15\Omega_{0}$ bottom.
@@
Figure 3: Longitudinally averaged profiles of the zonal wind at five different latitudes, spanning the region from the equator ($50^{\circ}$) to the northern pole ($0^{\circ}$), for three HD209458b-like models at different rotation rates: $0.125\Omega_{0}$ left, $\Omega_{0}$ centre, and $15\Omega_{0}$ right. These three rotation rates have been chosen to emphasise the change in zonal flow dynamics with rotation regime. Note: Positive wind speeds indicate eastwards flows.
@@ 3 Synchronous Rotation
We start by exploring how rotation affects the atmospheric dynamics of synchronously rotating exoplanets, focusing our analysis on three rotation rates ($0.125\Omega_{0}$, $\Omega_{0}$, $15\Omega_{0}$) which are typical of the three dynamical regions observed. We return to our full sample in subsection 3.5 where we investigate the influence of rotation on vertical heat transport.
@@ 3.1 Zonal-Mean Zonal-Winds
This narrowing of the jet and suppression of off-equator flows continues as we move to even higher rotation rates. However here another effect also becomes apparent: a reduction in the total zonal transport by the jet, which can be seen in 7(a) as a reduction in the mean zonal wind speed with increasing $Ro$ (an effect we discuss in more detail in subsection 3.4). As we discuss in subsection 3.4, this is likely due to the Coriolis suppression of the jet-driving off-equator standing Rossby and Kelvin waves.
@@
Figure 4: Zonally and temporally averaged meridional circulation streamfunctions for three HD209458b-like models with rotation rates of; $0.125\Omega_{0}$ top, $\Omega_{0}$ middle, and $15\Omega_{0}$ bottom. Note that the meridional circulation profile is plotted on a log scale with clockwise circulations shown in red and anti-clockwise circulations shown in blue. At the base rotation rate, for example, these circulations combine at the equator to drive an equatorial downflow at almost all pressures.
@@ 3.2 Meridional Circulation Streamfunction
Starting in the slowly rotating regime, we find (4(a)) a relatively weak meridional circulation profile that generally consists of a single circulation cell in each hemisphere. The only exception to this is the very deepest regions of the simulation domain which has not yet equilibrated due to the low local wind speeds (1(a)) and hence long dynamical time-scales. In the outer atmosphere, where the influence of rotation on the global overturning circulation is strongest, these circulations are driven by the net-eastward zonal wind, and as a result we find clockwise/anti-clockwise circulation cells in the southern/northern hemisphere respectively, cells that combine at the equator to drive a weak, net, downflow. However, as we move deeper ($P>0.5$\mathrm{bar}$$) and the net zonal-wind fades to near zero(see 3(a)), the sense of these circulations is reversed, leading to an equatorial upwelling that is typical of a global overturning circulation that is barely influence by rotation (Hammond & Lewis 2021).
@@
Figure 5: Temporally averaged zonal wind (arrows) and temperature profile (map) at three different pressures (0.0026 bar - top - to 0.2 bar - bottom) for three HD209458b-like models at different rotation rates ranging from $0.125\Omega_{0}$ left - to $15\Omega_{0}$ right.
3.3 Horizontal-Wind and Temperature Profiles
In order to better understand the dynamics discussed above, we next investigate the horizontal wind and temperature profiles at select pressure levels for three HD209458b-like models which fall into our three regimes of interest, as shown in Figure 5. Note that pressure levels have been chosen so as to emphasise the difference between the dynamical regimes (i.e. the differences associated with varying both rotation rate and relative radiative/advective dynamical strength).

### v2

@@ Abstract
Methods. We use DYNAMICO to run a series of long-timescale, HD209458b-like, atmospheric models at various rotation rates. For models that are tidally-locked, we consider rotation rates between $1/16^{\mathrm{th}}$ and $40$ times the rotation rate of HD209458b, whilst for non-synchronous models we consider the range $1/8^{\mathrm{th}}$ to $4$ times HD209458b.
Results. We find that our synchronous models fall into one of three $\Omega$-dependent regimes: at low $\Omega$, we find that the outer atmosphere dynamics are driven by a divergent day-night wind, which drives weak vertical transport and can lead to the formation of a night-side hot-spot. At intermediate $\Omega$, we find classical hot Jupiter dynamics, whilst at high $\Omega$ we find a strong Coriolis effect that suppresses off-equator dynamics, including the jet-driving standing waves, thus also reducing vertical transport. As for non-synchronicity, when small, like Showman et al. (2009), we find that it has little effect on the dynamics. However as it grows, we find that temporal variations prevents the formation of the persistent structures that drive large scale dynamics and transport.
@@ 1 Introduction
Thanks to the continuing efforts of the observational and instrumental astrophysics community, the parameter regime in which highly irradiated Jupiter-like planets can be found has expanded greatly (see, for example, Shporer et al. 2017; Hjorth et al. 2019; Huang et al. 2020; Schlecker et al. 2020; Dong et al. 2021; Stangret et al. 2022; Ulmer-Moll et al. 2022). In particular, there has been a significant increase in the range of orbital radii in which irradiated Jupiter-mass planets can be observed, leading to an increase in the number of known ‘warm’ Jupiters. Simultaneously, numerical studies of hot Jupiter, and hot Brown Dwarf, atmospheres has revealed that rotation can play a significant role on both the horizontal and vertical dynamics (Showman et al. 2015; Mendonça 2019; Sainsbury-Martinez et al. 2021). For example, Sainsbury-Martinez et al. (2021), found that, in addition to the link between radius inflation and planetary irradiation, a link might also exist between the level of radius inflation observed and the rotation rate of the exoplanet in question (see Figure 19). Specifically, they suggested that the lack of observed radius inflation for the ultra-short-orbit hot Brown Dwarf SDSS1411b could be explained by the influence that rotation has on the atmospheric dynamics. A similar dynamical story might also explain the radius of the ultra-short-orbit hot Brown Dwarf WD0137-349b, with the atmospheric models of Lee et al. (2020) also suggesting that the dynamics of this hot Brown Dwarf are highly influenced by rotation.
Examples of how the Coriolis effect can influence atmospheric dynamics can be found in our own solar system. Generally, planetary atmospheric dynamics depend upon the influence of both planetary irradiation and rotation, with the latter effect leading to the formation of three distinct atmospheric regimes.
At slow rotation rates, e.g. Venus and Titan, atmospheric dynamics are dominated by planetary scale Hadley cells (Hadley 1735; Schneider 2006; Showman et al. 2013b) which mix temperature horizontally, resulting in weak latitudinal temperature gradients. This is known as Cyclostrophic circulation, and leads to significant super-rotation of the winds, as well as strong, circumpolar, vortexes, the latter of which is described well by Thermal Wind Balance (Flasar et al. 2010). Note that, in the slow rotation regime, strong drag can also play a role in shaping the atmospheric wind, with a force balance between pressure-gradients and drag leading to winds to just flowing down pressure gradients rather than developing Cyclostrophic circulations (Showman et al. 2013a).
@@
The strength of the meridional heat transport continues to decrease as we move into the rapid rotation regime, as seen on Jupiter, Saturn, and the other gaseous solar-system giants. Here the Coriolis effect plays a significant role at all latitudes, leading to a Geostrophic atmosphere dominated by the infamous banded jet structure (Ingersoll & Vasavada 1998) (with baroclinic instabilities being proposed as a mechanism to drive mixing/angular-momentum-transport between bands; Young et al. 2019).
@@
A similar change in dynamics is expected to occur in exoplanetary atmospheres, albeit over a significantly wider parameter space thanks to the much wider variety of planetary type/orbital radii combinations found outside of our solar system. One such combination found only outside of our solar-system is the hot Jupiter, a Jupiter-like, gaseous planet which is in a very short orbit and hence is very highly irradiated. Whilst a number of recent studies have explored how atmospheric dynamics changes with orbital radii (Showman et al. 2015; Sainsbury-Martinez et al. 2023a), these come with the proviso that the models link effects caused by changes in both planetary irradiation and rotation, making it difficult to explore the effects that rotation alone has on the dynamics. Further, even studies which have explored the isolated effects of rotation alone (e.g; Mendonça 2019) have only done so for a few rotation rates, although even this limited study showed how rotation can shape the horizontal and vertical transport.
@@
As such, and in order to isolate how rotation affects the atmospheric dynamics (with a particular focus on vertical potential-temperature/enthalpy transport, which has the potential to explain the radius inflation problem for both hot Jupiters and hot Brown Dwarfs; Tremblin et al. 2017; Sainsbury-Martinez et al. 2019, 2021; Sainsbury-Martinez et al. 2023b), here we present a series of deep, HD209458b-like atmospheric models in which everything but the rotation rate is held constant. Specifically the rotation rate will be varied between $\Omega_{\mathrm{rot}}=0.0625\Omega_{0}$ and $\Omega_{\mathrm{rot}}=40\Omega_{0}$, where $\Omega_{0}=2.1\times 10^{-5}\mathrm{s^{-1}}$ is the canonical angular rotation rate of HD209458b, and the atmospheric models will be run until the deep atmosphere (with $P_{max}=200$\mathrm{bar}$$) has equilibrated. These values where chosen such that our models would span the rotation regimes found in our own solar system, and we will investigate cases in which the rotation is both synchronous and non-synchronous.
@@
The structure of this work is as follows: In section 2 we introduce the model we use to explore the effects of rotation on atmospheric dynamics and observables: This includes both a brief description of our 3D global circulation model (GCM) DYNAMICO, as well as a more indepth discussion of the parametrisations we use. Then, in section 3, we discuss how rotation influences the atmospheric dynamics of synchronously rotating hot Jupiters, starting with zonal-mean dynamics before moving onto the horizontal and vertical winds and heat transport, which we explore via both the Helmholtz wind decomposition as well as the enthalpy flux. Together these allow us to investigate how rotation influences deep atmospheric heating, and hence observed radius inflation. We then, in section 4, perform a similar analysis for a set of hot Jupiter atmospheric models in which the orbital period and planetary rotation period are allowed to diverge (i.e. non-synchronous rotation), investigating how sensitive the standing-wave driven dynamics, and hence the vertical heat transport, are to longitudinal shifts in the planetary irradiation profile. We finish, in section 5, with concluding remarks, discussing the implications of our results and suggestions for future studies, with next-generation GCMs, which will investigate the robustness of our results. In particular, we are very interested in confirming if the formation of a night-side hot-spot at slow rotation rates is just a feature of our atmospheric models, or if it is a robust feature that may have important implications for both the chemistry and observations of more slowly rotating hot Jupiters (i.e. hot Jupiters in longer orbits around massive/bright stars).
@@ 2.1 DYNAMICO
DYNAMICO is one of a number of highly computationally efficient global circulation models which has been developed to run on the next-generation of Exascale supercomputers, enabling atmospheric models of resolutions and regimes that where previously out of reach. To achieve this, DYNAMICO solves the primitive equations of meteorology (Dubos & Voitus 2014; Vallis 2006) on a spherical icosahedral grid (Dubos et al. 2015). Here we configure this horizontal-grid such that the angular resolution of each grid-cell is approximately $2.5^{\circ}$. Vertically, the levels of the pressure-grid are linearly spaced in $\log\left(P\right)$ space with a maximum pressure of $200$\mathrm{bar}$$ and a minimum pressure of $7\times 10^{-3}$\mathrm{bar}$$. This pressure range was selected such that we model both the highly irradiated outer atmosphere where irradiation always dominates over transport as well as a sufficiently large region of the deep atmosphere such that the effects of differing vertical heat transport rates are clearly visible. Note that the vertical boundaries of our model are closed and stress-free such that the only means of energy transport into and out-of the system are the numerical-dissipation (implemented as a hyperdiffusion filter with a fixed dissipation time-scale) required to stabilise the system against grid-scale noise, and the Newtonian cooling relaxation scheme which we use to model the stellar irradiation (subsection 2.2).
Following the calibration procedures of Sainsbury-Martinez et al. (2019) and Sainsbury-Martinez et al. (2021), we select a numerical dissipation time-scale of $\tau_{dissip}=2500\,\mathrm{s}$ in order to ensure the stability of the equatorially confined jets that develop in our most rapidly rotating models (e.g. 1(c)). These narrow jets with their high latitudinal velocity gradients also result in us setting a 90 second model time-step, which balances the need to correctly reproduce model physics, whilst also minimising the computational cost modelling equilibrium deep atmospheres.
@@ 2.2 Radiative forcing via Newtonian cooling
$\frac{H}{c_{p}}=\frac{\partial T\left(P,\theta,\phi\right)}{\partial t}=-\frac{T\left(P,\theta,\phi\right)-T_{eq}\left(P,\theta,\phi\right)}{\tau_{rad}\left(P\right)}\,,$ (1)
where $H$ is the local heating rate and $c_{p}$ is the specific heat capacity. For numerical reasons, we express the equilibrium temperature profile ($T_{eq}\left(P,\theta,\phi\right)$) onto which the atmosphere is relaxed, in terms of the night-side temperature profile $T_{n}\left(P\right)$ (Iro et al. 2005) and a pressure dependent day-night temperature difference ($\Delta{T}(P)$) which has been calibrated such as to reproduce observations of the hot Jupiter HD209458b:
@@
Note that, for the majority of models discussed here, we set the deep radiative time-scale to infinity $\tau_{rad}^{d}=\infty$, thus disabling radiative forcing in the deep atmosphere, between 20 and 200 bar, where it is expected to be significantly slower than advection/transport (Sainsbury-Martinez et al. 2019).
@@ 2.3 HD209458b-like Atmospheric Models
In order to explore how rotation impacts the outer and deep atmospheric dynamics and observables of highly irradiated, gaseous, exoplanets, we have explored synchronously rotating atmospheres at 28 different rotation rates (between $\Omega_{\mathrm{rot}}=0.0625\Omega_{0}$ and $\Omega_{\mathrm{rot}}=40\Omega_{0}$ discussed in section 3) and non-synchronously rotating atmospheres at 13 different planetary rotation rates (between $\Omega_{\mathrm{rot}}=0.25\Omega_{0}$ and $\Omega_{\mathrm{rot}}=4\Omega_{0}$, with the orbital period held constant at that of HD209458b - discussed in section 4). Here each model was run for over 200 Earth-years of simulation time to ensure that the radiatively active layers ($P<10$ bar) were equilibrated. Note that the deep atmospheres in all the models presented here should also be close to equilibrium, however the deep heating rate asymptotically slows as we approach steady-state and hence modelling the last few kelvin of heating would require significant additional computation for little change in the atmospheric dynamics (relative to the difference between a deep isotherm and adiabat; Sainsbury-Martinez et al. 2019).
@@
Figure 1: Zonally and temporally averaged, over five Earth-years of simulation time, zonal wind profiles for three HD209458b-like models with rotation rates of; $0.125\Omega_{0}$ top, $\Omega_{0}$ middle, and $15\Omega_{0}$ bottom.
@@
Figure 3: Longitudinally averaged profiles of the zonal wind at five different latitudes, spanning the region from the equator to the northern pole ($90^{\circ}$), for three HD209458b-like models at different rotation rates: $0.125\Omega_{0}$ left, $\Omega_{0}$ centre, and $15\Omega_{0}$ right. These three rotation rates have been chosen to emphasise the change in zonal flow dynamics with rotation regime. Note: Positive wind speeds indicate eastwards flows and for each wind profile we average over a $20^{\circ}$ latitudinal band centred at the labelled latitude.
@@ 3 Synchronous Rotation
We start by exploring how rotation affects the atmospheric dynamics of synchronously rotating exoplanets, focusing our analysis on three rotation rates ($0.125\Omega_{0}$, $\Omega_{0}$, $15\Omega_{0}$) which are typical of the three dynamical regions observed. We return to our full sample in subsection 3.5 where we investigate the influence of rotation on vertical heat transport. And we finish, in subsection 3.6, with a brief discussion of non-dimensional length scales, rotational regimes, and the resolution of our models.
@@ 3.1 Zonal-Mean Zonal-Winds
This narrowing of the jet and suppression of off-equator flows continues as we move to even higher rotation rates. However here another effect also becomes apparent: a reduction in the total zonal transport by the jet, which can be seen in 17(a) as a reduction in the mean zonal wind speed with increasing $Ro$ (an effect we discuss in more detail in subsection 3.4). As we discuss in subsection 3.4, this is likely due to the Coriolis suppression of the jet-driving off-equator standing Rossby and Kelvin waves.
@@
Figure 4: Zonally and temporally, over five Earth-years of simulation time, averaged meridional circulation streamfunctions for three HD209458b-like models with rotation rates of; $0.125\Omega_{0}$ top, $\Omega_{0}$ middle, and $15\Omega_{0}$ bottom. Note that the meridional circulation profile is plotted on a log scale with clockwise circulations shown in red and anti-clockwise circulations shown in blue. At the base rotation rate, for example, these circulations combine at the equator to drive an equatorial downflow at almost all pressures.
@@ 3.2 Meridional Circulation Streamfunction
Starting in the slowly rotating regime, we find (4(a)) a relatively weak meridional circulation profile that generally consists of a single circulation cell in each hemisphere. The only exception to this is the very deepest regions of the simulation domain which has not yet equilibrated due to the low local wind speeds (1(a)) and hence long dynamical time-scales. In the outer atmosphere, where the influence of rotation on the global overturning circulation is strongest, these circulations are driven by the net-eastward zonal wind, and as a result we find clockwise/anti-clockwise circulation cells in the southern/northern hemisphere respectively, cells that combine at the equator to drive a weak, net, downflow. However, as we move deeper ($P>0.5$\mathrm{bar}$$) and the net zonal-wind fades to near zero(see 3(a)), the sense of these circulations is reversed, leading to an equatorial upwelling that is typical of a global overturning circulation that is barely influenced by rotation (Hammond & Lewis 2021).
@@
Figure 5: Temporally averaged, over five Earth-years of simulation time, zonal wind (quivers) and temperature maps at three different pressures (0.0026 bar - top - to 0.2 bar - bottom) for three HD209458b-like models at different rotation rates ranging from $0.125\Omega_{0}$ left - to $15\Omega_{0}$ right. For each pressure level and rotation rate, the RMS wind speed is shown in the top right.
3.3 Horizontal-Wind and Temperature Maps
In order to better understand the dynamics discussed above, we next investigate the horizontal wind and temperature maps at select pressure levels for three HD209458b-like models which fall into our three regimes of interest, as shown in Figure 5. Note that pressure levels have been chosen so as to emphasise the difference between the dynamical regimes (i.e. the differences associated with varying both rotation rate and relative radiative/advective dynamical strength).